# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v1

@@ 4.1 Definitions and notations
A time series is a sequence of data points indexed in time order (that here is assumed to be regularly sampled, i.e., equispaced). An equispaced univariate time series of length $T$ is a sequence of scalar observations collected over time, denoted as $\vec{x}_{t}\in\mathbb{R}^{T}$. An equispaced multivariate time series of length $T$ is a sequence of $D$-dimensional vector observations collected over time, denoted as $\vec{x}_{t}\in\mathbb{R}^{D\times T}$.
@@
A graph $\mathcal{G}$ is a pair of finite sets $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ is a set of $n$ nodes (also called vertices) and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. In an undirected graph each edge is an unordered pair of nodes $\{v_{i},v_{j}\}$, while in a directed graph the edges have an orientation and correspond to ordered pairs $(v_{i},v_{j})$. When an edge $(v_{i},v_{j})$ or $\{v_{i},v_{j}\}$ exists, the nodes $v_{i}$ and $v_{j}$ are called adjacent.
@@ 4.2 Overview of graph neural networks
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, a GNN model (not necessarily changing in time, i.e., spatio-temporal) aims to generate an embedding (i.e., a vector of real entries) for each node in the graph. An overall graph embedding consists of representing a graph into a possibly different-dimensional space while preserving its structural information.
The notion of GNN was initially proposed by Gori et al. in 2005 in Ref. [gori_aggiunto] and by Scarselli et al. in 2009 in Ref. [Scarselli2009] as an architecture to implement a function that maps a graph $\mathcal{G}$ with $n$ nodes into an $m$-dimensional Euclidean space. In Ref. [Scarselli2009] the target node’s embedding was learned by first propagating forward at fixed weights neighbor information within the graph in an iterative manner, until a stable fixed point is reached (whose existence is guaranteed under certain assumptions by Banach’s fixed point theorem), and then propagating errors on targets back across the graph. The calculation of node embeddings is in general quite complex, and over time many other methods have been developed. These methods are based on different architectures and sophisticated mechanisms. Based on these aggregating mechanisms, a taxonomy of GNN models with different classes has been defined. In this review, we adopt a taxonomy for the discussed spatio-temporal GNNs similar to the one proposed by Chen et al. in Ref. [survey_Lucheroni]. More details will be given in Subsec. 4.3.
@@ 4.3 Spatio-temporal GNNs and their taxonomy
The majority of research on spatio-temporal GNNs is focused on multivariate time series, as they can be naturally abstracted into spatio-temporal graphs, as depicted in Fig. 5. In this context, nodes correspond to different variables at a given time step, and edges represent relationships between variables.
@@
This spatio-temporal modeling assumes that the feature information of a node depends on both its own historical values and on the historical data from its neighboring nodes [survey_spatiotemporal]. Once the the graph structure is given, information is propagated within it until the output is obtained, as better explained in the following.
When it comes to the development of a GNN model, there are two primary approaches: one that handles spatial and temporal substructures in separate blocks, and another that integrates and processes these two substructures together [survey_spatiotemporal]. This is implemented by means of modules that are purely spatial, purely temporal, or hybrid combinations of the two. These modules are then organized into a series of blocks to create the final spatio-temporal GNN model.
In the taxonomy of spatio-temporal GNNs used in this review, the focus is exclusively on the description of the spatial module that propagates the information within the nodes at fixed time $t$ (not explicitly included in the following equations for simplicity), which is in the reviewed literature a fundamental aspect that differentiates one GNN model from another. Based on the mechanism at the basis of the information propagation in the graph at a given time step, spatio-temporal GNNs can be further classified as recurrent GNNs, convolutional GNNs, and attentional GNNs, as explained in the following. These classes are not mutually exclusive, as some models may fit into hybrid categories. In addition, the classes are not fully exhaustive, as there may be models that do not align with any of the existing categories. However, for the purpose of this review, the taxonomy discussed above is sufficient.
Recurrent GNNs. Recurrent graph neural network modeling mainly includes the results of early studies on GNNs, where the node embeddings are generated by iteratively propagating neighbor information until a stable fixed point is reached. The iterative equation for the calculation of the embedding (or hidden state) of node $u$ has the form
@@
where $k$ is the index related to iteration, $f$ is a parametric trainable function, $x_{u}$ is the attribute associated to node $u$, $r_{(u,v)}$ is the weight of the edge $(u,v)$ between the nodes $u$ and $v$, $\mathcal{N}(v)$ denotes the neighborhood of node $v$, and the initial hidden states $h^{(0)}$ are usually randomly initialized. As detailed in Ref. [Scarselli2009], which introduced recurrent GNNs, to ensure the convergence of Eq. (2), it is necessary that the iterative function $f$ is a contraction mapping.
Convolutional GNNs. Convolutional graph neural networks are closely related to recurrent graph neural networks. They were at first introduced by Kipf and Welling in 2017 in Ref. [graph_convolution_kipf]. Unlike recurrent GNNs, which apply the same iterative contraction mapping $f$ until an equilibrium is reached, convolutional GNNs use different parameters at each updating step, by stacking multiple graph convolutional layers to extract the node embeddings [cit_rev3].
A basic form for the equation for the $k^{\text{th}}$ layer of a convolutional graph neural network with $n$ nodes is
$H^{(k)}=f_{k}\left(\tilde{\mathcal{A}}H^{(k-1)}\Theta_{k-1}\right)\>,\quad\tilde{\mathcal{A}}=\hat{\mathcal{D}}^{-\frac{1}{2}}\hat{\mathcal{A}}\hat{\mathcal{D}}^{-\frac{1}{2}}\>,\quad\hat{\mathcal{A}}=\mathcal{A}+I\>,\quad\hat{\mathcal{D}}_{ii}=\sum_{j}\hat{\mathcal{A}}_{ij}\>,\quad\quad H^{(0)}=X\>,$ (3)
where $f_{k}$ represents the convolutional operation performed at layer $k\in\{1,\dots,K\}$, $\mathcal{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, $I\in\mathbb{R}^{n\times n}$ is the identity matrix (used to add the self-loops in the adjacency matrix $\mathcal{A}$), $\hat{\mathcal{D}}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix associated to $\hat{\mathcal{A}}$, $H^{(k)}\in\mathbb{R}^{n\times d_{k}}$ are the $d_{k}$-dimensional node embeddings $h^{(k)}_{v}$ of the entire graph produced by the $k^{\text{th}}$ layer, $X\in\mathbb{R}^{n\times d_{0}}$ is the collection of all node attributes $x_{v}$ of the entire graph, and $\Theta_{k}\in\mathbb{R}^{d_{k-1}\times d_{k}}$ is a trainable weight matrix for the $k^{\text{th}}$ layer [survey_Lucheroni]. The mechanism of node embedding generation from Eq. (3) is illustrated in Fig. 6(b), where it is put in contrast with the mechanism of a recurrent GNNs, shown in Fig. 6(a).
@@
Eq. (4a) represents a learnable linear transformation of the node embedding $h_{u}^{(k)}$ using the trainable weight matrix $W_{k}$ of layer $k$. Eq. (4b) computes a pairwise masked attention score between nodes $u$ and $v$ by concatenating (||||| |) the vectors $z_{u}^{(k)}$ and $z_{v}^{(k)}$, calculating a dot product with a learnable weight vector $\vec{a}^{\,(k)}$, and applying the LeakyReLU nonlinearity. The quantity $\epsilon_{uv}$ indicates the importance of the features of node $v$ to node $u$. The term ”masked” refers to the fact that $\epsilon_{uv}$ are only computed for neighboring nodes. Eq. (4c) normalizes the attention score through the softmax function in order to make the $\alpha_{uv}$ comparable across different nodes. Finally, Eq. (4d) computes the next layer embedding $h_{u}^{(k+1)}$ by aggregating the information in neighboring nodes, weighted by the attention scores [GAT]. Graph attention networks can also further use a multi-head attention mechanism to stabilize the learning process of self-attention. In practice, the node updating operation in graph attention networks is a generalization of the traditional averaging or max-pooling of neighboring nodes, which allows each node to compute a weighted average of its neighbors and identify the most relevant ones [Brody2021HowAA].
As anticipated, the taxonomy discussed so far focuses only on the propagation mechanisms of the spatial component. However, spatio-temporal GNNs typically consist of a stack of spatial, temporal, or hybrid modules, each one with distinct roles.
The spatial module is responsible for propagating information between nodes, thereby enabling the analysis of cross-sectional inter-dependencies between different variables. As previously explained in the description of the mechanisms underlying the taxonomy, the spatial module aggregates information starting from immediate neighboring nodes and extends it outward, thereby capturing also the influence of distant nodes. Noticeably, this spatial operation does not account in principle for temporal information. In contrast, the temporal module, often based on architectures such as LSTM, GRU, or self-attention, focuses on the evolution of data over time, independently from cross-sectional node interactions.
When these different modules are stacked together, spatial and temporal information end up to be used simultaneously, although they are processed in separate modules. This occurs because the computation of each node state is influenced by both spatial and temporal information. This combined approach allows the spatio-temporal GNN model to effectively capture at the same time both spatial relationships and temporal dynamics.
@@ 4.4 Determination of the graph structure
A crucial issue when dealing with spatio-temporal GNNs for time series problems is the determination of the graph structure, i.e. the connectivity of the nodes. Namely, some time series datasets come by nature with a pre-defined graph structure (e.g., road network), while some others do not. When a natural pre-defined graph structure is available, it helps the model to better capture the underlying dynamics of the system. When it is not directly available, it must be in some way defined by the user (based on domain knowledge or some metrics), determined by some algorithm (such as visibility graphs [visibility_Gori], [visibility_graph], [art_14]), or learned by the model itself.
@@
Once the graph structure has been determined, it is necessary to define the weighted graph adjacency matrix, which is a generalization of the graph adjacency matrix whose entries are given by the edge weights of the graph. The edge weights can be defined a priori by the user, or again, learned by the model itself based on a pre-defined architecture. In the first case, the edge weights are assigned based on some pre-defined metrics or criteria chosen by the user (e.g., spatial distance between the sites or similarity measures). In the latter case, the weights are continuously adjusted throughout training as the model learns from the data.
Each section of the following thematic analysis examines which approach is prevalent.

### v2

@@ 4.1 Definitions and notations
A time series is a sequence of data points indexed by timestamps, that here are assumed to be equispaced. An equispaced univariate time series of length $T$ is a sequence of scalar observations collected over time, denoted as $\vec{x}_{t}\in\mathbb{R}^{T}$. An equispaced multivariate time series of length $T$ is a sequence of $D$-dimensional vector observations collected over time, denoted as $\vec{x}_{t}\in\mathbb{R}^{D\times T}$.
@@
A graph $\mathcal{G}$ is a pair of finite sets $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ is a set of $n$ nodes (also called vertices) and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of connecting edges. In an undirected graph each edge is an unordered pair of nodes $\{v_{i},v_{j}\}$, while in a directed graph the edges have an orientation, and correspond to ordered pairs $(v_{i},v_{j})$. When an edge $(v_{i},v_{j})$ or $\{v_{i},v_{j}\}$ exists, the nodes $v_{i}$ and $v_{j}$ are called adjacent.
@@
Hereafter, the time index will be dropped to indicate graphs not necessarily changing in time, i.e., not necessarily spatio-temporal.
@@ 4.2 Overview of graph neural networks
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, a GNN model (not necessarily spatio-temporal) aims to generate an embedding (i.e., a vector of real entries) for each node in the graph. An overall graph embedding consists of representing a graph in a space of chosen dimension, while preserving its structural information.
The notion of GNN was initially proposed by Gori et al. in 2005 in Ref. [gori_aggiunto] and by Scarselli et al. in 2009 in Ref. [Scarselli2009] as an architecture designed to implement a function that maps a graph $\mathcal{G}$ with $n$ nodes into an $m$-dimensional Euclidean space. Noticeably, this definition was intended in general for graph data, not specifically time-dependent data, which is the focus of this SLR. In Ref. [Scarselli2009] the target node’s embedding was learned in two steps. First, neighbor information is iteratively propagated with fixed weights within the graph until a stable fixed point is reached (whose existence is guaranteed under certain assumptions by Banach’s fixed point theorem). Second, errors on targets are propagated back across the graph. Notice that in this setup, propagation is performed in “internal time”, not in “clock time”. More in general, the calculation of node embeddings is quite complex, and over time many other methods have been developed. These methods are based on different architectures and sophisticated aggregation mechanisms. Based on these aggregating mechanisms, a taxonomy of GNN models with different classes has been defined. In this review, we adopt a taxonomy for the discussed spatio-temporal GNNs similar to the one proposed by Chen et al. in Ref. [survey_Lucheroni]. More details about that will be given in Subsec. 4.3.
@@ 4.3 Spatio-temporal GNNs and their taxonomy
A large part of the research on spatio-temporal GNNs is focused on multivariate time series, as they can be naturally abstracted into spatio-temporal graphs, as depicted in Fig. 5. In this context, nodes usually correspond to different variables at a given time step, and edges represent relationships between variables.
@@
This specific form of spatio-temporal modeling assumes that the feature information of a node depends on both its own historical values and on the historical data from its neighboring nodes [survey_spatiotemporal]. Notice that once the the graph structure is given, information is propagated (in “internal time”) within it until the output is obtained, as better explained in the following. Our discussion will build up from this perspective.
When it comes to the development of a spatio-temporal GNN model, there are two primary approaches: one that handles spatial and temporal substructures in separate blocks, and another that integrates and processes these two substructures at once [survey_spatiotemporal]. These approaches are implemented by means of modules that are purely spatial, purely temporal, or hybrid combinations of the two. These modules can then be organized into a series of blocks, thus creating the sought spatio-temporal GNN model.
In the taxonomy of spatio-temporal GNNs used in this review, the focus is exclusively on the description of the spatial module that propagates the information within the nodes at fixed time $t$ (not explicitly included in the following equations for simplicity), which, in the reviewed literature, is a fundamental aspect that differentiates one GNN model from another. Based on the mechanism at the basis of the information propagation in the graph at a given time step, spatio-temporal GNNs can be further classified as recurrent GNNs, convolutional GNNs, and attentional GNNs, as discussed in the following. These classes are not mutually exclusive, as some models may fit into hybrid categories. In addition, these classes are not fully exhaustive of the variety of available models, as there may be models that do not align with any of the existing categories. However, for the purpose of this review, the taxonomy discussed above will be considered sufficient.
Recurrent GNNs. Recurrent GNN modeling mainly proceeds from the results of early studies on GNNs, where the node embeddings are generated by iteratively propagating neighbor information until a stable fixed point is reached. The iterative equation for the calculation of the embedding (or hidden state) of node $u$ has the form
@@
where $k$ is the index related to iteration (hence running on “internal time”), $f$ is a parametric trainable function, $x_{u}$ is the attribute associated to node $u$, $r_{(u,v)}$ is the weight of the edge $(u,v)$ between the nodes $u$ and $v$, $\mathcal{N}(v)$ denotes the neighborhood of node $v$, and the initial hidden states $h^{(0)}$ are usually randomly initialized. As detailed in the seminal Ref. [Scarselli2009], which introduced recurrent GNNs, in order to ensure the convergence of Eq. (2) it is necessary that the iterative function $f$ is a contraction mapping.
Convolutional GNNs. Convolutional GNNs are closely related to recurrent GNNs. They were introduced by Kipf and Welling in 2017 in Ref. [graph_convolution_kipf]. Unlike recurrent GNNs, which apply the same iterative contraction mapping $f$ until an equilibrium is reached, convolutional GNNs can use different parameters at each updating step, by stacking multiple graph convolutional layers to extract the node embeddings [cit_rev3]. Noticeably, as the number of convolutional layers increases, the convolutional perceptual field expands, and a more abstract representation of information is thus obtained [art_461].
A basic form for the equation for the $k^{\text{th}}$ layer of a convolutional GNN with $n$ nodes is
$\begin{gathered}H^{(k)}=f_{k}\left(\tilde{\mathcal{A}}H^{(k-1)}\Theta_{k-1}\right)\>,\quad\tilde{\mathcal{A}}=\hat{\mathcal{D}}^{-\frac{1}{2}}\hat{\mathcal{A}}\hat{\mathcal{D}}^{-\frac{1}{2}}\>,\quad\hat{\mathcal{A}}=\mathcal{A}+I\>,\quad\\ \hat{\mathcal{D}}_{ii}=\sum_{j}\hat{\mathcal{A}}_{ij}\>,\quad\quad H^{(0)}=X\>,\end{gathered}$ (3)
where $f_{k}$ represents the convolutional operation performed at layer $k\in\{1,\dots,K\}$ (with $K$ typically kept small to prevent issues like over-smoothing [oversmoothing]), $\mathcal{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, $I\in\mathbb{R}^{n\times n}$ is the identity matrix (used to add self-loops in the adjacency matrix $\mathcal{A}$), $\hat{\mathcal{D}}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix associated to $\hat{\mathcal{A}}$, $H^{(k)}\in\mathbb{R}^{n\times d_{k}}$ are the $d_{k}$-dimensional node embeddings $h^{(k)}_{v}$ of the entire graph produced by the $k^{\text{th}}$ layer, $X\in\mathbb{R}^{n\times d_{0}}$ is the collection of all node attributes $x_{v}$ of the entire graph, and $\Theta_{k}\in\mathbb{R}^{d_{k-1}\times d_{k}}$ is a trainable weight matrix for the $k^{\text{th}}$ layer [survey_Lucheroni]. The mechanism of node embedding generation from Eq. (3) is illustrated in Fig. 6(b), where it is put in contrast with the mechanism of a recurrent GNNs, shown in Fig. 6(a).
@@
Eq. (4a) represents a learnable linear transformation of the node embedding $h_{u}^{(k)}$ using the trainable weight matrix $W_{k}$ of layer $k$. Eq. (4b) computes a pairwise masked attention score between nodes $u$ and $v$. This is done by applying the concatenation operation ||||| | to the vectors $z_{u}^{(k)}$ and $z_{v}^{(k)}$, then calculating a dot product of the concatenation with a learnable weight vector $\vec{a}^{\,(k)}$, and finally applying the LeakyReLU nonlinearity. The quantity $\epsilon_{uv}$ indicates the importance of the features of node $v$ to node $u$. The term “masked” refers to the fact that $\epsilon_{uv}$ are computed for neighboring nodes only. Eq. (4c) normalizes the attention score through the softmax function in order to make the $\alpha_{uv}$ comparable across different nodes. Finally, Eq. (4d) computes the embedding $h_{u}^{(k+1)}$ of the next layer by aggregating the information in neighboring nodes, weighted by the attention scores [GAT]. Graph attention networks can in addition use a multi-head attention mechanism to stabilize the learning process of self-attention. In practice, the node updating operation in graph attention networks is a generalization of the traditional averaging or max-pooling of neighboring nodes, which allows each node to compute a weighted average of its neighbors and identify the most relevant ones [Brody2021HowAA].
As anticipated, the taxonomy discussed so far focuses only on the propagation mechanisms of the spatial component. However, spatio-temporal GNNs typically consist of a stack of spatial, temporal, or also hybrid modules, each one with distinct roles. The function of these modules is detailed hereafter.
The spatial module is responsible for propagating information between nodes, thereby enabling the analysis of cross-sectional inter-dependencies between different variables. As previously discussed in the description of the mechanisms underlying the assumed taxonomy, the spatial module aggregates information starting from immediate neighboring nodes and extends it outward, thereby capturing also the influence of distant nodes. This spatial operation cannot directly account for temporal information. In contrast, the temporal module, often based on architectures such as LSTM, GRU, or self-attention, focuses on the evolution of data over time, independently from cross-sectional node interactions [art_384]. When these different modules are stacked together, spatial and temporal information end up to be used simultaneously, although they are processed in separate modules. This occurs because the computation of each node state is influenced by both spatial and temporal information. This combined approach allows the spatio-temporal GNN model to be a powerful architecture able to effectively capture at the same time both spatial relationships and temporal dynamics.
@@ 4.4 Determination of the graph structure
A crucial issue when dealing with spatio-temporal GNNs for time series problems is the determination of the graph structure, i.e., the connectivity of the nodes. Namely, some time series datasets come by their own nature with a pre-defined graph structure (e.g., a road network), whereas some others do not. When a natural pre-defined graph structure is available, it helps the model to better capture the underlying dynamics of the system. When it is not directly available, it must be defined by the user in some way (based on domain knowledge or some metrics), or determined by some algorithm (such as visibility graphs [visibility_Gori, visibility_graph, art_14]), or learned by the model itself.
@@
Once the graph structure is determined, it is necessary to define the weighted graph adjacency matrix, which is a generalization of the graph adjacency matrix $\mathcal{A}$ whose entries are given by the edge weights of the graph. The edge weights can be defined a priori by the user, or again, learned by the model itself based on a pre-defined architecture. In the first case, the edge weights are assigned based on some pre-defined metrics or criteria chosen by the user, for example by the spatial distance between the sites, or by similarity measures. In the latter case, the weights are continuously adjusted throughout training as the model keeps learning from the data.
In the following, we present a concise overview of the most common techniques for defining graph structures.
Pre-defined graphs are typically generated based on prior knowledge and specific metrics. These methods for the construction of the graphs typically rely on distance or similarity information. As for location-based techniques, they utilize geographical coordinates (latitude and the longitude) of the nodes to calculate the distance between them (e.g., using haversine or geodesic distance). This distance information is then used to determine the edge weights with techniques such as the thresholded Gaussian kernel (which applies a negative exponential function to the distance), weighted $k$-NN, or reversed MinMax. As for similarity-based techniques instead, they rely on similarity (or dissimilarity) of signal data. These techniques include methods such as correlation, cosine similarity, dynamic time warping (DTW), maximum information coefficient (MIM), or transfer entropy. In general, no single approach stands out as a clear winner, and performance should be evaluated on each dataset. For references and a more detailed explanation of these and other techniques for constructing pre-defined structures, see Ref. [art_160].
As for automatic graph structure learning techniques, a distinction can be made between models that rely on predefined graph structures (sometimes referred to as supplementary graph structure learning models), and those that learn the graph structure entirely from scratch. In the first category, there are models such as Graph WaveNet (GWN) [GWN], which constructs an adaptive matrix as a complement to a predefined graph structure [art_405]. These techniques still rely on some form of prior assumptions. To address this limitation, models that learn the graph structure from scratch have been proposed. One of them is the MTGNN model [MTGNN], which extracts uni-directional relationships from node embeddings as it learns, and achieves state-of-the-art results across different datasets. It is one of the first models to use self-learned graph structures [art_417]. When the graph structure is learned, some authors introduce a graph regularization term to the training loss function to optimize the learned graph adjacency matrix. This regularization term is typically designed to control the sparsity of the matrix, and to promote graph proximity, encouraging larger entries when the series in the corresponding nodes are “close enough” [art_448]. Some researchers argue that models which generate graphs without prior knowledge may overlook local conditions when constructing adjacency matrices [art_397]. Therefore, it is important to conduct further research in this area to identify the most effective techniques for generating informative graphs. For a more detailed explanation of graph learning models, see Refs. [art_396, art_405].
Another key aspect in the determination of the graph structure is the choice between using a static or dynamic adjacency matrix, i.e. whether to use the same matrix for all graphs or not. When working with graphs defined in advance by specific metrics, it is common to use either a static matrix or a dynamic matrix generated through a sliding window approach. In contrast, when graphs are learned by the model, dynamic adjacency matrices are typically employed. The use of a dynamic adjacency matrix has the advantage of capturing the dynamic nature of the problem [art_447]. However, the learned dynamic graphs often show significant differences between adjacent time steps, whereas real-world graph structures typically evolve smoothly over time. Therefore, this approach may not accurately capture the stable evolution observed in the real world.
Consequently, some researchers are exploring an intermediate approach, assuming that adjacent graph structures follow a basic pattern. Instead of defining static or dynamic adjacency matrices, they fuse a static backbone graph with a dynamic adjacency matrix. The static network captures the common graph structure over time, while the dynamic network focuses on learning the changes over time. An example of this intermediate approach can be found in Ref. [art_415], which also provides a comparison between this intermediate approach and the classical ones.
In general, there is no rule that can determine in advance the best approach to follow. In fact, it depends on both the application domain and the specific dataset. For this reason, we do not explore this aspect further. Each subsection of the thematic analysis in Sec. 5 discusses the most prevalent approach for each group of papers.
4.5 Model evaluation
Evaluating the performance of a time series forecasting or classification model is essential to assess its reliability. Moreover, experimental results often show that even when datasets involve the same forecasting task, the transferability of model architectures is limited [art_408]. Therefore, a proper evaluation of the model is a crucial aspect. A first key aspect to consider is its accuracy. The accuracy of a model can be measured by many different metrics, which vary between classification and forecasting tasks. Each metric provides its own insight into different aspects of prediction quality, so multiple metrics are typically used for a more comprehensive evaluation. In addition to presenting a table of the error metrics computed for the entire graph, some authors provide a spatial visualization of these errors across the various nodes, as for example in Refs. [art_28, art_449].
As for forecasting problems, commonly used evaluation metrics include the mean absolute error (MAE), mean squared error (MSE), root mean squared error (RMSE), mean absolute percentage error (MAPE), root relative squared error (RRSE), the coefficient of determination (R2), and the correlation coefficient (CORR). Let $N$ denote the number of observations, $y_{i}$ the true value associated with observation $i$, $\hat{y}_{i}$ the corresponding output of the model, and the bar denote the average. The formulas for these metrics (for a single-step forecasting horizon) are provided in Tab. 5.
Table 5: Evaluation metrics for single-horizon forecasting tasks and their formulas.
Metrics Formula
Mean absolute error (MAE) $\frac{1}{N}\sum_{i=1}^{N}|y_{i}-\hat{y}_{i}|$
Mean squared error (MSE) $\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}$
Root mean squared error (RMSE) $\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}$
Mean absolute percentage error (MAPE) $\frac{1}{N}\sum_{i=1}^{N}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|\cdot 100$
Root relative squared error (RRSE) $\sqrt{\frac{\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}}$
Coefficient of determination (R2) $1-\frac{\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}$
Correlation coefficient (CORR) $\frac{\sum_{i=1}^{N}(y_{i}-\bar{y})(\hat{y}_{i}-\bar{\hat{y}})}{\sqrt{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}\sum_{i=1}^{N}(\hat{y}_{i}-\bar{\hat{y}})^{2}}}$
Similar formulas can be applied to multi-horizon forecasting problems. All the metrics listed, except for the last one, are considered better when their values are lower, as they indicate smaller discrepancies between the predicted and actual values. In contrast, CORR quantifies the linear correlation between actual values and forecasts, and a high correlation is preferable.
As for classification problems, common evaluation metrics include accuracy, precision, recall, and F1 score. Given $TP$, $TN$, $FP$, and $FN$ the number of true positives, true negatives, false positives, and false negatives, respectively, the formulas for the binary classification problem are reported in Tab. 6.
Table 6: Evaluation metrics for binary classification tasks and their formulas.
Metrics Formula
Accuracy $\frac{TP+TN}{TP+TN+FP+FN}$
Precision $\frac{TP}{TP+FP}$
Recall $\frac{TP}{TP+FN}$
F1 score $\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
Similar formulas can be applied for the multi-class classification problem.
Besides accuracy, another important aspect to consider is the complexity of the model. Finding an optimal balance between accuracy and complexity is crucial for developing models that are not only accurate but also efficient. For this reason, some authors include the training and inference computational times, as well as the number of parameters of their models, allowing for a comparison of these aspects across different models. However, not all authors do so, and even if they do it, the characteristics of the hardware and other settings used by each author must be taken into account when comparing computational times. Consequently, it was not possible to provide an accurate and detailed comparison of the relationship between complexity and performance across the different models in this review. Although creating a plot to compare accuracy and computational times of the different models would be very useful (as in Ref. [art_407]), it is not possible to provide such a comparison here. The authors of the various papers should be encouraged to incorporate such comparisons in their work.